A charge Q1=-28.9 μC is held stationary at the origin. How much energy is required to bring another charge q2=-96.6 μC from "infinity" (that is from a very very large distance from each other) to the location (2.5 cm,1.0 cm)?

The answer should have correct sign and be in appropriate units.
Answer #1

given,

Q1=-28.9 μC

q2=-96.6 μC

distance = sqrt(2.5^2 + 1^2)

distance = 2.6926 cm

electric potnetial energy = kq1 * q2 / r

electric potnetial energy = 9 * 10^9 * -28.9 * 10^-6 * -96.6 * 10^-6 / 0.026926

energy required = 933.137 J